# Flexible MOF Generation with Torsion-Aware Flow Matching

### v1

@@ F.1 Training details
Appendix G Property evaluation for structure prediction
Figure 8 presents the property distributions for the structure prediction task (Section 6.1). All baselines closely align with the ground-truth, indicating that key MOF properties are well preserved.
Figure 8: Property distributions. We compare MOF property distributions of the ground-truth, DiffCSP, MOFFlow, and MOFFlow-2. The distribution has been smoothed with kernel density estimation. All baselines closely follow the ground-truth distribution.
Appendix H Additional related work
Flow matching. Flow matching (FM) is a simulation-free method for training continuous normalizing flows [CNF; 50] by learning a time-dependent vector field that transports samples from a source distribution to a target distribution [27, 51, 52]. Since its introduction, FM has been extended in several directions, such as generalizations to Riemannian manifolds [26] and formulations based on optimal transport [53, 54]. These advances facilitated its application to protein design [55, 56, 57, 22], molecular conformation generation [58], and materials generation [9, 10].
Generative models for inorganic crystals. Most crystal generation models to date have focused on inorganic crystals. CDVAE [59] introduces a latent diffusion model with SE(3)-equivariant encoders and decoders to preserve periodicity. DiffCSP [8] jointly denoises the lattice and coordinates information using a periodic E(3)-equivariant architecture. EquiCSP [60] accounts for lattice permutation symmetry, while DiffCSP++ [61] incorporates space group constraints. FlowMM [9] replaces diffusion with an equivariant flow-based model, improving sampling efficiency. MatterGen [18] enables property-guided generation through conditional diffusion. More recently, ADiT [24] proposes a unified Transformer architecture for modeling both molecular and crystalline systems.

### v3

@@ F.1 Training details
Appendix G Additional experiments
G.1 Property evaluation for structure prediction
Figure˜8 presents the property distributions for the structure prediction task (Section˜6.1). All baselines closely align with the ground-truth, indicating that key MOF properties are well preserved.
Figure 8: Key property distributions of MOFs generated by DiffCSP, MOFFlow, and MOFFlow-2. Distributions are smoothed using kernel density estimation. All baseline methods closely match the reference data distribution.
[…]
(b) Energy per atom of generated MOFs
[…]
G.3 Energy-level evaluation using MLIP
We further evaluate the energy levels of generated MOFs using UMA [52], a state-of-the-art machine-learned interatomic potential. Specifically, we generate 10,000 structures with both MOFFlow-2 and MOFDiff and compute the energy per atom (eV/atom) using the uma-m-1.1 model. As shown in Figure˜9(b), the energy distribution of structures generated by MOFFlow-2 aligns more closely with the training dataset than that of MOFDiff.
Appendix H Limitations
Although MOFFlow-2 demonstrates strong potential for MOF structure prediction and generation, several limitations remain. Firstly, because the pipeline relies on RDKit for initial conformer generation, it is challenging to predict structures whose organic building blocks are chemically invalid or incomplete; for example, those lacking carboxylate groups, which are common in MOF decomposition schemes [37]. Second, for conditional generation in Section˜G.2, only the building block generator is conditioned on the property, whereas the structure prediction module may also depend on it. Finally, our evaluation of energy levels with machine learning interatomic potentials is less accurate than that with density functional theory (DFT).